### v1

# Cash Flow Underwriting with Bank Transaction Data: Advancing MSME Financial Inclusion in Malaysia

Abstract.
Despite accounting for 96.1% of all businesses in Malaysia (Department of Statistics Malaysia, 2025), access to financing remains one of the most persistent challenges faced by Micro, Small, and Medium Enterprises (MSMEs). Newly established or young businesses are often excluded from formal credit markets as traditional underwriting approaches rely heavily on credit bureau data. This study investigates the potential of bank statement data as an alternative data source for credit assessment to promote financial inclusion in emerging markets. Firstly, we propose a cash flow-based underwriting pipeline where we utilise bank statement data for end to end data extraction and machine learning credit scoring. Secondly, we introduce a novel dataset of 611 loan applicants from a Malaysian lending institution. Thirdly, we develop and evaluate credit scoring models based on application information and bank transaction-derived features. Empirical results show that the use of such data boosts the performance of all models on our dataset, which can improve credit scoring for new-to-lending MSMEs. Lastly, we intend to release the anonymised bank transaction dataset to facilitate further research on MSMEs financial inclusion within Malaysia’s emerging economy.
Bank Statement Data, Alternative Data, Credit Scoring
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on AI in Finance; November 15–18, 2025; Singaporeccs: Computing methodologies Classification and regression trees
Figure 1. Proposed end-to-end cash flow underwriting workflow leveraging bank statement transaction data for credit scoring.:
1. Introduction
Financial inclusion remains a pressing challenge in emerging markets, where a significant portion of the population lacks sufficient credit history to access formal lending. According to the Securities Commission Malaysia (Malaysia, 2025), MSMEs are estimated to contribute around 60% of Malaysia’s gross domestic product. Despite their significant contributions, MSMEs remain underserved by financial institutions in terms of access to financing, giving rise to an estimated MYR 90 billion funding gap (Baharin, 2024; Corporation, 2017).
One fundamental issue is that many MSMEs do not have lending history. Traditional credit assessment relies heavily on credit bureau data such as repayment history, outstanding obligations, and past delinquencies. While this model works well for established businesses, it has shortcomings for MSMEs with thin credit files:
(i) It is inherently backward-looking and focuses on past repayment behaviour rather than current or forward-looking capacity to repay.
(ii) It overlooks real-time financial signals or operational dynamics that more accurately reflect a firm’s present financial health.
(iii) It omits alternative indicators of creditworthiness, such as cash flow consistency, receivables and payables patterns, digital transactions, and behavioural metrics from daily financial activity.
On the other hand, bank statements represent an up-to-date and verifiable source of financial behaviour which also capture income regularity, spending patterns, and cash flow stability. This study explores the use of Malaysian bank statement transactions as alternative data for credit scoring model development and evaluates its predictive value for MSMEs. The objectives are as follows:
  • To propose a cash flow-based underwriting workflow capable of ingesting and analysing bank transaction data for credit decisioning to narrow the MSMEs financing gap.
  • To introduce the first ever Malaysian bank statement dataset based on loan application of MSMEs.
  • To evaluate the performance of machine learning based credit scoring models trained on the proposed dataset to study the feasibility and predictive power of transaction-based features in assessing the creditworthiness of MSMEs.
2. Related Works
With the recent technological advancements in machine learning algorithms, the research community begins to realize the limitations of existing traditional credit scoring methods (Gote and Mendhe, 2024). Such studies are important to enhance financial inclusion of MSMEs in emerging markets where MSMEs are often the backbone of economic growth although they have limited access to financing (Elebe and Imediegwu, 2021). As such, alternative data from non-traditional sources that are not included in standard credit bureau files are being used for credit scoring (Group, 2024).
2.1. Limitations of Traditional Credit Scoring
Traditional credit scoring models used by the financial institutions primarily rely on credit bureau data such as payment history, amount of debt, and other indicators (Shivhare, 2024). Unfortunately, reliance on such data creates a high entry barrier for MSMEs in emerging economies where they usually lack audited financial statements or loan servicing history required by financial institutions (Courchane and Baines, 2020; Elebe and Imediegwu, 2021). This leads MSMEs to be perceived as high risk applicants and results in a perpetual cycle of financial exclusion that severely limits their growth potential (Courchane and Baines, 2020). Besides, traditional methods fail to capture recent cash flow status of MSMEs, which can be an accurate representation of financial health (Elebe and Imediegwu, 2021). Recent survey indicates that 58% of lending institutions feel less confident to make decisions based on traditional credit data only (Watterson, 2024). The increasing dissatisfaction motivates the urgent need to use alternative data for a more adaptive and inclusive credit assessment (Gote and Mendhe, 2024).
2.2. Credit Scoring with Alternative Data
In order to overcome the limitations of traditional methods, credit underwriting using alternative data has emerged as a new approach (Gote and Mendhe, 2024; Watterson, 2024; Ngwenya, 2024). For instance, mobile network data are used for credit scoring in Africa (Ngwenya, 2024; Gathu, 2020) and bank account transactions are used for cash flow underwriting in consumer lending (Watterson, 2024). Recent studies suggest that incorporating transactional data can improve predictive accuracy and expand credit to underserved populations (Djeundje et al., 2021). For instance, research has shown that retail transaction data can help construct alternative credit scores (Lee et al., ). A study examining two Indian financial technology (FinTech) firms found that transactional data from platform activity worked as well as credit bureau data in predicting creditworthiness and improved the predictive accuracy when combined with bureau data (Caire and Vidal, 2024). However, the adoption of alternative data in MSMEs lending in Malaysia remains at an early stage, with limited research focus from the financial industry.
2.3. Machine Learning Models in Credit Scoring
Machine learning (ML) models have proven to be highly effective in processing financial data to produce accurate credit scores (Bücker et al., 2022; Gunnarsson et al., 2021; Trivedi, 2020). These include statistical models like Logistic Regression, Naive Bayes (NB) and ensemble methods like Random Forest, AdaBoost, and Gradient Boosting (Bücker et al., 2022; Trivedi, 2020). For example, a Random Forest model achieved the best performance on Taiwanese credit card data (Abbas and Hussein, 2024). A NB classifier performed well on a highly imbalanced dataset from a New Zealand lender, especially when undeclared features derived from bank statements were incorporated to supplement application form data (Bunker et al., 2016). Given the positive impact of ML models in credit scoring (Bello, 2024), we aim to improve financial inclusion of Malaysian MSMEs through our ML-based cash flow underwriting workflow.
3. Bank Statement Cash Flow Underwriting
The proposed end-to-end cash flow underwriting workflow enables financial institutions to integrate bank statement transaction data analysis into their credit decision-making process. Its primary objective is to enhance traditional credit underwriting by leveraging bank statement-derived features. As illustrated in Figure 1, the workflow consists of three main layers:
(i) Web Layer (Customer Onboarding): This layer acts as the entry point for all loan application-related data including bank statements. A web-based platform is built to allow MSME owners to submit loan applications and for bank officers to review them.
(ii) Application Layer (Bank Statement Analyser): This layer processes uploaded bank statements from the web interface. Multiple specialized engines automate the extraction and conversion of unstructured transactions into structured data, analyses cash flow-based indicators and bank statement related features, and applies rule-based checks to detect potential fraud and data abnormalities.
(iii) Data & Scoring Layer (Cash Flow Underwriting): This layer stores the analysed data and engineered features in a secure database to ensure traceability, regulatory compliance, and auditability. The stored data are then preprocessed, and feature selection methods are applied to retain the most informative variables. Using the selected features, a predictive model is trained to estimate the probability of default and classify credit risk.
The proposed workflow offers several advantages for financial institutions. It shortens turnaround time by automating manual tasks, thereby improving operational efficiency. It also expands credit access for thin-file MSMEs by leveraging transactional data instead of relying primarily on historical credit bureau records, allowing MSMEs with limited or no bureau data to obtain financing.
4. Bank Statement Transaction Dataset
Table 1. Statistics of MSME loan application datasets.
Split Non-Event Event Total
Train (60%) 310 56 367
Validation (40%) 208 37 245
Overall 518 93 611
Transaction data derived from bank statements represents a valuable alternative resource for cash flow underwriting. To the best of our knowledge, no published studies have examined the use of bank statement data for credit assessment of MSMEs in Malaysia. In collaboration with a lending institution, we constructed the first Malaysian bank statements dataset to address this gap.
This study adopts the Cross-Industry Standard Process for Data Mining (CRISP–DM) framework (Chapman et al., 2000) as the methodological basis. CRISP–DM is a widely recognized process model for systematic FinTech studies (Cheng, 2023; Rawat, 2023; da Rocha and de Sousa Junior, 2010) and comprises six main phases i.e. business understanding, data understanding, data preparation, modeling, evaluation, and deployment. The first phase involves evaluating bank-statement transaction data as an alternative source for MSME credit risk assessment in Malaysia. The second phase focuses on constructing the proposed dataset of 611 MSME loan applicants and is split into training and validation sets with a 60:40 ratio as shown in Table 1. Among these, 518 applicants have a good credit record, while 93 have a history of default. Each applicant’s record contains two main components: application form information submitted during the loan application process (e.g., demographic and business characteristics) and bank statement transaction data capturing detailed inflows and outflows over a six-month period. The third phase is about data preparation and includes several key tasks:
(i) Data cleaning and de-duplication to ensure consistency.
(ii) Fixing missing values and standardizing transaction categories.
(iii) Deriving features and variables from transaction data, such as determining cash flow stability, deposit regularity, and balance volatility. All personally identifiable information was anonymised prior to analysis. To preserve confidentiality, the features are grouped into application form information (7 features) and bank statement features (10 features). Although the specific feature calculations cannot be disclosed due to a non-disclosure agreement, all bank statement features are derived solely from bank statement data.
5. Transaction Data-based Credit Scoring
The fourth phase of CRISP–DM in this study employs Logistic Regression (LR) (Cox, 1958) on application form and bank transaction data derived from our dataset as the baseline credit scoring model. LR is widely adopted in credit risk modeling due to its interpretability, statistical robustness, and capacity to estimate the probability of default (Bücker et al., 2022; Trivedi, 2020; Abbas and Hussein, 2024; Bunker et al., 2016). To ensure that only informative predictors are retained, feature selection is guided by Information Value (IV), a widely used metric in credit risk modeling for quantifying the predictive power of individual variables (Ngwenya, 2024). IV measures how well a feature distinguishes between default and non-default outcomes, with higher values indicating stronger discriminatory capacity. Following industry practice (Siddiqi, 2017), we adopt the commonly used interpretive thresholds: $\mathrm{IV}<0.02$ (not predictive), $0.02\leq\mathrm{IV}<0.1$ (weak), $0.1\leq\mathrm{IV}<0.3$ (medium), $0.3\leq\mathrm{IV}<0.5$ (strong), and $\mathrm{IV}\geq 0.5$ (suspiciously high; suggesting potential data leakage).
Figure 2. ROC curves of all models evaluated on the validation split of the proposed dataset.:
(a) IV of features derived from application and bank statement data.:
(b) AUROC on train and validation splits with different combinations.:
Figure 3. Results of ablation studies on IV and AUROC performance on the proposed dataset. Best viewed in colour.
5.1. Logistic Regression Model
Formally, let $\mathbf{x}_{i}\in\mathbb{R}^{d}$ denote the feature vector for applicant $i$, where $d$ is the number of features engineered from both application form data and bank statement transactions, and let $y_{i}\in\{0,1\}$ be the binary output indicating default ($y_{i}=1$) or non-default ($y_{i}=0$). The LR model specifies the conditional probability of default as
(1) $P(y_{i}=1\mid\mathbf{x}_{i};\boldsymbol{\beta})=\sigma(\beta_{0}+\mathbf{x}_{i}^{\top}\boldsymbol{\beta}),$
where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function, $\beta_{0}\in\mathbb{R}$ is the intercept term, and $\boldsymbol{\beta}\in\mathbb{R}^{d}$ is the coefficient vector associated with the predictors. The parameters are estimated by maximizing the penalized log-likelihood function:
(2) $\mathcal{L}(\boldsymbol{\beta})=\sum_{i=1}^{n}\left[y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i})\right]-\lambda\,\|\boldsymbol{\beta}\|_{2}^{2},$
where $p_{i}=P(y_{i}=1\mid\mathbf{x}_{i};\boldsymbol{\beta})$ and $\lambda\geq 0$ controls the strength of the $\ell_{2}$ regularization (Ng, 2004). This mitigates overfitting by shrinking coefficient magnitudes, which is particularly important when modeling high-dimensional, transaction-derived features.
6. Experiments
Aligned with the fifth phase of CRISP–DM, we conduct experiments to evaluate the proposed cash flow underwriting workflow and the role of bank statement data in MSME credit scoring. The experiments include comparisons with established machine learning methods and ablation studies on transaction-derived features.
6.1. Implementation Details
We benchmark the baseline LR model against several widely used machine learning methods, including Random Forest (RF) (Breiman, 2001), Gradient Boosting (GB) (Friedman, 2001), and AdaBoost (AB) (Schapire, 2013). All models are implemented in scikit-learn (scikit learn, 2025c, a, b, d). For the results in Section 6.2, models were trained with default hyperparameters. In Section 6.3, hyperparameters were tuned using a randomized grid search with 50 trials to ensure robustness. Model performance is evaluated using the Area Under the Receiver Operating Characteristic Curve (AUROC) (Bradley, 1997), which measures the ability to discriminate between default and non-default cases across varying thresholds. An AUROC of 0.5 indicates no discriminative power (equivalent to random guessing), whereas a value of 1.0 indicates perfect discrimination.
6.2. Quantitative Results
Quantitative results on the validation split of the proposed dataset are shown in Figure 2. The LR model achieves the highest AUROC of 0.782, representing a 30.77% improvement over AB, the lowest-performing model with a score of 0.598. RF follows with 0.655, and GB with 0.633. This ranking can be explained by the fact that LR is well suited to capturing the linear structure of the data, whereas more complex ensemble methods such as RF, GB, and AB are prone to overfitting when applied to relatively small samples, particularly under class imbalance (310 non-event vs. 56 event cases in the training set). In such settings, models that provide well-calibrated probabilities and strong generalization tend to perform better, which is consistent with prior benchmarking studies showing that LR often matches or outperforms tree-based ensembles on small-to-medium credit datasets (Baesens et al., 2003; Lessmann et al., 2015; Brown and Mues, 2012).
6.3. Ablation Studies
Results in Figure 3(a) show that bank statement–derived features exhibit stronger predictive power than application form information. Nine of the ten transaction-based features rank above all application features, with the sole exception of customer classification, a pre-assigned label from the lending institution describing the business entity. Customer classification slightly surpasses the recent three-month maximum average balance with a relative difference of 7.92%, while also maintaining a substantial margin of 8.67% over the next highest-ranked application feature. These findings reinforce the superior discriminatory strength of bank statement–derived variables in distinguishing default from non-default outcomes.
Furthermore, we conducted ablation experiments by progressively removing feature groups to study which features matter most for MSME credit scoring. We applied 5-fold cross-validation on the training split of the proposed dataset to reduce the risk that results are due to random chance and to increase confidence in model stability. From results illustrated in Figure 3(b), we observe a substantial uplift when incorporating bank transaction data. The AUROC of LR improves from 0.672 (application information only) to 0.821 (bank transaction only), representing an uplift of 0.149 points (approximately 22% relative improvement). When both feature sets are combined, the AUROC of LR further increases to 0.8497, confirming that transactional data provides significant incremental predictive value. The validation results are consistent with the cross-validation findings. On both splits, all models that use bank transaction data outperform those based solely on application information. The combined features lead to the best result for all models. In general, these findings strongly support the hypothesis that bank statement transaction data offers significant predictive power in credit risk assessment for MSMEs in Malaysia. While application form information provides some discriminatory power, the inclusion of bank transaction-based features substantially enhances model performance, reinforcing the effectiveness of the proposed cash flow underwriting workflow.
7. Conclusion and Future Work
In conclusion, this study introduces bank statement transaction data as alternative data for MSMEs loan underwriting in Malaysia. Our study demonstrates that transaction-derived features capture dynamic aspects of MSMEs’ financial behaviour that are often overlooked in traditional credit models. These features also exhibit strong predictive power with high IV scores. Empirical results show that models built solely on application information perform modestly, whereas those based on bank transaction data achieve substantially better results. The combined features further improve the model’s predictive power, confirming the complementary nature of alternative data. In the future, we aim to integrate real-time transactional data into our workflow to better assess the creditworthiness of MSMEs with limited credit history in Malaysia.