### v3

# Cash Flow Underwriting with Bank Transaction Data: Advancing MSME Financial Inclusion in Malaysia

Abstract
Despite accounting for 96.1% of all businesses in Malaysia (smeprofile), access to financing remains one of the most persistent challenges faced by Micro, Small, and Medium Enterprises (MSMEs). Newly established businesses are often excluded from formal credit markets as traditional underwriting approaches rely heavily on credit bureau data. This study investigates the potential of bank statement data as an alternative data source for credit assessment to promote financial inclusion in emerging markets. First, we propose a cash flow-based underwriting pipeline where we utilise bank statement data for end-to-end data extraction and machine learning credit scoring. Second, we introduce a novel dataset of 611 loan applicants from a Malaysian lending institution. Third, we develop and evaluate credit scoring models based on application information and bank transaction-derived features. Empirical results show that the use of such data boosts the performance of all models on our dataset, which can improve credit scoring for new-to-lending MSMEs. Finally, we will release the anonymised bank transaction dataset to facilitate further research on MSME financial inclusion within Malaysia’s emerging economy.
1 Introduction
Figure 1: Proposed end-to-end cash flow underwriting workflow leveraging bank statement transaction data for credit scoring.
Financial inclusion remains a pressing challenge in emerging markets, where a significant portion of the population lacks sufficient credit history to access formal lending. According to the Securities Commission Malaysia (cm_plan), MSMEs are estimated to contribute around 60% of Malaysia’s gross domestic product. Despite their significant contributions, MSMEs remain underserved by financial institutions in terms of access to financing, giving rise to an estimated MYR 90 billion funding gap (international2017msme).
One fundamental issue is that many MSMEs do not have lending history. Traditional credit assessment relies heavily on credit bureau data such as repayment history, outstanding obligations, and past delinquencies. While this traditional credit model works well for established businesses, it has shortcomings for MSMEs with thin credit files:
  1. i
    It is inherently backward-looking and focuses on past repayment behaviour rather than current or forward-looking capacity to repay.
  2. $i$
    It overlooks real-time financial signals or operational dynamics that could more accurately reflect a firm’s present financial health.
  3. iii
    It omits alternative indicators of creditworthiness, such as cash flow consistency, receivables and payables patterns, digital transactions, and behavioural metrics from daily financial activity.
On the other hand, bank statements represent an up-to-date and verifiable source of financial behaviour which also capture income regularity, spending patterns, and cash flow stability. This study explores the use of Malaysian bank statement transactions as alternative data for credit scoring model development and evaluates its predictive value for MSMEs. The objectives are as follows:
  • To propose a cash flow-based underwriting workflow capable of ingesting and analysing bank transaction data for credit decisioning to narrow the MSMEs financing gap.
  • To introduce the first ever Malaysian bank statement dataset based on loan applications of MSMEs.
  • To evaluate the performance of machine learning–based credit scoring models trained on the proposed bank statement transaction dataset, and to examine the feasibility and predictive power of transaction-derived features in assessing MSME creditworthiness.
2 Related Work
With the recent technological advancements in machine learning algorithms, the research community has recognised the limitations of existing traditional credit scoring methods (gote2024building). Such studies are crucial for enhancing financial inclusion of MSMEs in emerging markets, where MSMEs are often the backbone of economic growth despite having limited access to financing (elebe2021credit). As such, alternative data from non-traditional sources that are not included in standard credit bureau files are being used for credit scoring (world_bank).
Limitations of Traditional Credit Scoring
Traditional credit scoring models used by financial institutions primarily rely on credit bureau data such as payment history, amount of debt, and other indicators (shivhare2024utilize). Unfortunately, reliance on such data creates a high entry barrier for MSMEs in emerging economies where they usually lack audited financial statements or loan servicing history required by financial institutions (courchane2019use; elebe2021credit). This leads MSMEs to be perceived as high-risk applicants and results in a perpetual cycle of financial exclusion that severely limits their growth potential (courchane2019use). Besides, traditional methods fail to capture recent cash flow status of MSMEs, which can be an accurate representation of financial health (elebe2021credit). A recent survey indicates that 58% of lending institutions feel less confident to make decisions based on traditional credit data only (datos). The increasing dissatisfaction motivates the urgent need to use alternative data for a more adaptive and inclusive credit assessment (gote2024building).
Credit Scoring with Alternative Data
In order to overcome the limitations of traditional methods, credit underwriting using alternative data has emerged as a new approach (gote2024building; datos; africa). For instance, mobile network data are used for credit scoring in Africa (africa; kenya) and bank account transactions are used for cash flow underwriting in consumer lending (datos). Recent studies suggest that incorporating transactional data can improve predictive accuracy and expand credit to underserved populations (djeundje2021enhancing). For instance, research has shown that retail transaction data can help construct alternative credit scores (express_peru). A study examining two Indian financial technology (FinTech) firms found that transactional data from platform activity worked as well as credit bureau data in predicting creditworthiness and improved the predictive accuracy when combined with bureau data (cgap). However, the adoption of alternative data in MSME lending in Malaysia remains at an early stage, with limited research focus from the industry.
Machine Learning Models in Credit Scoring
Machine learning (ML) models have proven to be highly effective in processing financial data to produce accurate credit scores (bucker2022transparency; gunnarsson2021deep; trivedi2020study). These include statistical models like Logistic Regression, Naive Bayes (NB), and ensemble methods like Random Forest, AdaBoost, and Gradient Boosting (bucker2022transparency; trivedi2020study). For example, a Random Forest model achieved the best performance on Taiwanese credit card data (abbas2024algorithm). A NB classifier performed well on a highly imbalanced dataset from a New Zealand lender, especially when undeclared features derived from bank statements were incorporated to supplement application form data (bunker2016improving). Given the positive impact of ML models in credit scoring, we aim to improve financial inclusion of Malaysian MSMEs through our ML-based cash flow underwriting workflow.
Figure 2: Overview of AI modules with their objectives and example outputs within the Bank Statement Analyser.
Agentic Workflow in Credit Scoring
The deployment of Large Language Model (LLM)-powered agentic systems marks a major advancement in automated credit scoring, loan approval, and customer risk profiling within financial services (okpala2025agentic; ali2025efficient; paleti2024agentic). Recent agentic frameworks have demonstrated strong performance in complex model risk management tasks, including credit card approval and portfolio risk modelling (okpala2025agentic). Beyond these use cases, agentic workflows are transforming banking by enabling adaptive credit profiling and predictive loan approvals, thereby improving risk assessment accuracy through bias detection and intelligent decision automation (paleti2024agentic). Similar approaches have been explored in mortgage lending, where machine learning–driven agents streamline underwriting and accelerate credit evaluation (chitturi2025agentic). However, existing methods remain limited in production deployment and lack a fully grounded, end-to-end workflow for MSMEs credit scoring.
3 Bank Statement Cash Flow Underwriting
The proposed end-to-end cash flow underwriting workflow enables financial institutions to integrate bank statement transaction data analysis into their credit decision-making process. Its primary objective is to enhance traditional credit underwriting by leveraging bank statement-derived features. As illustrated in Figure 1, the workflow consists of three main layers, which serve different purposes.
Web Layer (Customer Onboarding)
This layer acts as the entry point for all loan application-related data including bank statements. A web-based platform is built to allow MSME owners to submit loan applications and for bank officers to review them.
Application Layer (Bank Statement Analyser)
This layer processes uploaded bank statements from the web interface. Multiple specialised engines automate the extraction and conversion of unstructured transactions into structured data, analyse cash flow-based indicators and bank statement related features, and apply rule-based checks to detect potential fraud and data abnormalities. Within these extraction and analytic engines, several Artificial Intelligence (AI) modules perform complex and dedicated tasks to derive actionable insights from the uploaded bank statements as depicted in Figure 2. The outputs of these modules form the foundation for downstream credit scoring.
Key Information Extraction Module
The key information extraction module employs Optical Character Recognition (OCR) to extract essential fields from bank statements, including the account number, account holder name, address, and statement date. These attributes are cross validated across statements to ensure accurate document ownership verification. These fields enable the engine to link multiple accounts belonging to the same business entity, providing a holistic view of its financial activity across accounts.
Transaction Table Extraction Module
The transaction table extraction module localizes and digitizes tabular transaction data using OCR and layout analysis. It identifies table headers, then maps rows and columns to extract transaction date, description, debit, credit, and balance amounts. Additional pre-processing handles merged cells, page breaks, and format inconsistencies to preserve data fidelity. The resulting structured transaction records serve as critical inputs for downstream analytical modules that evaluate transaction patterns and account-level cash flow dynamics.
Fraud Analysis Module
The fraud analysis module leverages Computer Vision and rule-based techniques to detect tampered or falsified bank statements. For each document, it examines both visual and metadata cues such as inconsistent font styles or sizes, altered layouts, metadata mismatches, and abnormal pixel-level patterns that may indicate digital editing. Once potential tampering is identified, the affected statements are automatically flagged and returned to bank officers for manual verification. This module provides an additional layer of security and assurance within the underwriting workflow, ensuring that only authentic bank statements are analysed for credit scoring.
Network Analysis Module
The network analysis module constructs transaction networks from the extracted transaction data to uncover relationships among senders and beneficiaries across multiple bank accounts. Using graph-based algorithms, it maps transaction linkages to detect circular fund flows, kiting activities, and transfers to blacklisted or interrelated entities. This module also quantifies network connectivity and transaction density to reveal abnormal money movement patterns that may indicate fraudulent activity or financial distress, thereby enhancing overall risk assessment.
Cash Flow Analysis Module
The cash flow analysis module evaluates cash flow status based on the inflows and outflows extracted from transaction records. It then applies Natural Language Processing to infer transaction intent and classify entries into meaningful categories based on their descriptions. Using these classifications, this module groups related transactions to identify spending and receipt patterns, enabling the detection of outliers that deviate from typical cash flow behaviour. It also computes key metrics such as average, highest, and lowest balances, quantifying cash flow health and provides predictive features for credit scoring.
Data & Scoring Layer (Cash Flow Underwriting)
This layer stores the analysed data and engineered features in a secure database to ensure traceability, regulatory compliance, and auditability. The stored data are then preprocessed, and feature selection methods are applied to retain the most informative variables. Using the selected features, a predictive model is trained to estimate the probability of default and classify credit risk.
The proposed workflow offers several advantages for financial institutions. It shortens turnaround time by automating manual tasks, thereby improving operational efficiency. It also expands credit access for thin-file MSMEs by leveraging transactional data instead of relying primarily on historical credit bureau records, allowing MSMEs with limited or no bureau data to obtain financing.
4 Bank Statement Transaction Dataset
Table 1: Statistics of MSME loan application datasets.
Split Non-Event Event Total
Train (60%) 310 56 367
Validation (40%) 208 37 245
Overall 518 93 611
Transaction data derived from bank statements represents a valuable alternative resource for cash flow underwriting. To the best of our knowledge, no published studies have examined the use of bank statement data for credit assessment of MSMEs in Malaysia. In collaboration with a lending institution, we constructed the first Malaysian bank statements dataset to address this gap.
This study adopts the Cross-Industry Standard Process for Data Mining (CRISP–DM) framework (crisp_dm) as the methodological basis. CRISP–DM is a widely recognised process model for systematic FinTech studies (cheng2023evaluating; rawat2023applying; da2010identifying) and comprises six main phases i.e. business understanding, data understanding, data preparation, modelling, evaluation, and deployment. The first phase involves evaluating bank statement transaction data as an alternative source for MSME credit risk assessment in Malaysia. The second phase focuses on constructing the proposed dataset of 611 MSME loan applicants and is split into training and validation sets with a 60:40 ratio as shown in Table 1. Among these, 518 applicants have a good credit record, while 93 have a history of default. Each applicant’s record contains two main components: application form information submitted during the loan application process (e.g., demographic and business characteristics) and bank statement transaction data capturing detailed inflows and outflows over a six-month period. The third phase involves data preparation and includes several key tasks:
  1. i
    Cleaning and de-duplicating data to ensure consistency.
  2. $i$
    Fixing missing values, standardising transaction categories, and validating data integrity across all records.
  3. iii
    Deriving features and variables from transaction data, such as determining cash flow stability, deposit regularity, and balance volatility.
All personally identifiable information was anonymised prior to analysis. To preserve confidentiality, the features are grouped into application form information (7 features) and bank statement features (10 features). Although the specific feature calculations cannot be disclosed due to a non-disclosure agreement, all bank statement features are derived solely from bank statement data.
5 Transaction Data-based Credit Scoring
The fourth phase of CRISP–DM in this study employs Logistic Regression (LR) (cox1958regression) on application form and bank transaction data derived from our dataset as the baseline credit scoring model. LR is widely adopted in credit risk modelling due to its interpretability, statistical robustness, and capacity to estimate the probability of default (bucker2022transparency; trivedi2020study; abbas2024algorithm; bunker2016improving). To ensure that only informative predictors are retained, feature selection and transformation are guided by the Weight of Evidence (WOE) and Information Value (IV) framework, which are widely used in credit risk modelling for quantifying the predictive power of individual variables (africa). In practice, each feature is divided into discrete intervals based on suitable thresholds determined from the data. These bins allow the calculation of WOE and IV at the group level, which makes them easier to interpret within a logistic regression model. The following notation is used to quantify the distribution of defaults and non-defaults across feature bins:
Let $y_{i}\in\{0,1\}$ indicate default ($y_{i}=1$) or non-default ($y_{i}=0$) for applicant $i$. The total number of default and non-default applicants is defined as:
$N_{b}$ $=\sum_{i=1}^{n}\mathbb{I}(y_{i}=1),$ $N_{g}$ $=\sum_{i=1}^{n}\mathbb{I}(y_{i}=0),$ (1)
Furthermore, let $x_{ij}$ denote the value of feature $j$ for applicant $i$. For a given feature $j$, suppose it is divided into $K_{j}$ disjoint bins $\{B_{j1},\dots,B_{jK_{j}}\}$, then the corresponding counts within each bin is defined as:
$n_{bjk}$ $=\sum_{i=1}^{n}\mathbb{I}(x_{ij}\in B_{jk})\,\mathbb{I}(y_{i}=1),$
$n_{gjk}$ $=\sum_{i=1}^{n}\mathbb{I}(x_{ij}\in B_{jk})\,\mathbb{I}(y_{i}=0).$ (2)
The distribution of defaults and non-defaults in each bin are:
$\text{Dist}^{(b)}_{jk}$ $=\frac{n_{bjk}}{N_{b}},$ $\text{Dist}^{(g)}_{jk}$ $=\frac{n_{gjk}}{N_{g}}.$ (3)
Weight of Evidence (WOE)
Firstly, we calculate the WOE for bin $k$ of feature $j$ using:
$\mathrm{WOE}_{jk}=\log\!\left(\frac{\text{Dist}^{(g)}_{jk}}{\text{Dist}^{(b)}_{jk}}\right)=\log\!\left(\frac{n_{gjk}/N_{g}}{n_{bjk}/N_{b}}\right).$ (4)
A positive $\mathrm{WOE}_{jk}$ indicates that bin $B_{jk}$ is more common among non-default cases, suggesting lower risk, whereas a negative value indicates a higher likelihood of default. Once WOE is computed for each bin, the overall predictive strength of feature $j$ can then be summarised by its IV.
Information Value (IV)
Secondly, we measure the IV of feature $j$ by aggregating the WOE across all respective bins:
$\mathrm{IV}_{j}=\sum_{k=1}^{K_{j}}\left(\text{Dist}^{(g)}_{jk}-\text{Dist}^{(b)}_{jk}\right)\mathrm{WOE}_{jk}.$ (5)
A larger $\mathrm{IV}_{j}$ indicates stronger discriminatory power of feature $j$ between default and non-default classes. In this study, IV serves as both a feature-ranking criterion and an interpretability measure, helping to identify the most influential transaction-level indicators for creditworthiness. Following industry practice (intel_credit_scoring), we adopt the commonly used interpretive thresholds: $\mathrm{IV}<0.02$ (not predictive), $0.02\leq\mathrm{IV}<0.1$ (weak), $0.1\leq\mathrm{IV}<0.3$ (medium), $0.3\leq\mathrm{IV}<0.5$ (strong), and $\mathrm{IV}\geq 0.5$ (suspiciously high; suggesting potential data leakage).
Binning of Bank Statement Features
Specifically, we use quantile or supervised monotonic binning to obtain $\{B_{jk}\}_{k=1}^{K_{j}}$ for continuous transaction-derived variables (e.g., cash flow stability, balance volatility). Then, we group rare datapoints to ensure each bin has sufficient defaults and non-defaults for sparse categorical variables (e.g., business sector code). Computations of WOE and IV are performed on training folds only to avoid data leakage.
Logistic Regression Model
Given that WOE provides a log-odds transformation of each feature, this aligns naturally with the LR model. The WOE transformed feature value can be represented as $\mathrm{WOE}_{j}(x_{ij})$ and the LR model can be expressed in a simplified form as:
$\log\frac{P(y_{i}=1\,|\,\mathbf{x}_{i})}{P(y_{i}=0\,|\,\mathbf{x}_{i})}=\beta_{0}+\sum_{j=1}^{d}\beta_{j}\,\mathrm{WOE}_{j}(x_{ij}).$ (6)
where $\beta_{0}$ is the intercept and $\beta_{j}$ represents the coefficient for feature $j$. This allows direct interpretation of coefficients in terms of the credit risk associated with each feature. Positive $\beta_{j}$ values indicate that higher WOE corresponds to lower default risk, and vice versa. Hence, WOE encoding stabilises estimation and supports consistent, monotonic relationships between predictors and credit outcomes.
Formally, let $\mathbf{x}_{i}\in\mathbb{R}^{d}$ denote the feature vector for applicant $i$, where $d$ is the number of features engineered from both application form data and bank statement transactions, and let $y_{i}\in\{0,1\}$ be the binary output indicating default ($y_{i}=1$) or non-default ($y_{i}=0$). The LR model specifies the conditional probability of default as
$P(y_{i}=1\mid\mathbf{x}_{i};\boldsymbol{\beta})=\sigma(\beta_{0}+\mathbf{x}_{i}^{\top}\boldsymbol{\beta}),$ (7)
where $\sigma(z)=\frac{1}{1+e^{-z}}$ is the sigmoid function, $\beta_{0}\in\mathbb{R}$ is the intercept term, and $\boldsymbol{\beta}\in\mathbb{R}^{d}$ is the coefficient vector associated with the predictors. The parameters are estimated by maximizing the penalized log-likelihood function:
$\mathcal{L}(\boldsymbol{\beta})=\sum_{i=1}^{n}\left[y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i})\right]-\lambda\,\|\boldsymbol{\beta}\|_{2}^{2},$ (8)
where $p_{i}=P(y_{i}=1\mid\mathbf{x}_{i};\boldsymbol{\beta})$ and $\lambda\geq 0$ controls the strength of the $\ell_{2}$ regularization (ng2004feature). This mitigates overfitting by shrinking coefficient magnitudes, which is particularly important when modelling high-dimensional, transaction-derived features.
6 Experiments and Deployment
Following the fifth phase of CRISP–DM, we conduct a series of experiments to evaluate the effectiveness of the proposed cash flow underwriting workflow and the role of bank statement data in enhancing MSME credit scoring performance. The experiments include comparative analyses with established machine learning methods and detailed ablation studies on transaction-derived features.
Figure 3: ROC curves of all models evaluated on the validation split of the proposed dataset.
(a) IV of features derived from application and bank statement data.
(b) AUROC on train and validation splits with different combinations.
Figure 4: Results of ablation studies on IV and AUROC performance on the proposed dataset. Best viewed in colour.
Implementation Details
We benchmark the baseline LR model against several widely used machine learning methods, including Random Forest (RF) (breiman2001random), Gradient Boosting (GB) (friedman2001greedy), and AdaBoost (AB) (schapire2013explaining). All models are implemented in scikit-learn (sklearn_lr; sklearn_ada; sklearn_gb; sklearn_rf). For the results in Section 6, models were trained with default hyperparameters. In Section 6, hyperparameters were tuned using a randomized grid search with 50 trials to ensure robustness. We evaluate model performance using the Area Under the Receiver Operating Characteristic Curve (AUROC) (bradley1997use), which measures the ability to discriminate between default and non-default cases across varying thresholds. An AUROC of 0.5 indicates no discriminative power (equivalent to random guessing), whereas a value of 1.0 indicates perfect discrimination. To ensure interpretability and systematic analysis, both application and bank statement features are grouped into three categories:
  1. i
    Account Behaviour: logarithmic growth rate of average balances, average balance over the past six months, ratio between lowest balance in recent three months, percentage difference between the lowest balance, and highest average balance for the past three months.
  2. $i$
    Repayment Capacity: repayment capability calculated based on monthly installment.
  3. iii
    Business Demographics: years in business, business location, business sector code, number of directors, and director’s minimum age.
This structured feature grouping facilitates clearer attribution of model performance to different behavioural and demographic dimensions of MSME credit profiles. It also supports experimental reproducibility, while respecting confidentiality constraints around proprietary feature derivations.
Quantitative Results
Quantitative results on the validation split of the proposed dataset are shown in Figure 3. The LR model achieves the highest AUROC of 0.782, representing a 30.77% improvement over AB, the lowest-performing model with a score of 0.598. RF follows with 0.655, and GB with 0.633. This ranking can be explained by the fact that LR is well suited to capturing the linear structure of the data, whereas more complex ensemble methods such as RF, GB, and AB are prone to overfitting when applied to relatively small samples, particularly under class imbalance (310 non-event vs. 56 event cases in the training set). In such settings, models that provide well-calibrated probabilities and strong generalisation tend to perform better, which is consistent with prior benchmarking studies showing that LR often matches or outperforms tree-based ensembles on small-to-medium credit datasets (brown2012experimental; lessmann2015benchmarking).
Ablation Studies
Results in Figure 4(a) show that bank statement–derived features exhibit stronger predictive power than application form information. Nine of the ten transaction-based features rank above all application features, with the sole exception of customer classification, a pre-assigned label from the lending institution describing the business entity. Customer classification slightly surpasses the recent three-month maximum average balance with a relative difference of 7.92%, while also maintaining a substantial margin of 8.67% over the next highest-ranked application feature. These findings reinforce the superior discriminatory strength of bank statement features in distinguishing default from non-default outcomes.
In addition, we conducted ablation experiments by progressively removing feature groups to study which features matter most for MSMEs credit scoring. We applied 5-fold cross-validation on the training split of the proposed dataset to reduce the risk that results are due to random chance and to increase confidence in model stability. From results illustrated in Figure 4(b), we observe a substantial uplift when incorporating bank transaction data. The AUROC of LR improves from 0.672 (application information only) to 0.821 (bank transaction only), representing an uplift of 0.149 points (approximately 22% relative improvement). When both feature sets are combined, the AUROC of LR further increases to 0.8497, confirming that transactional data provides significant incremental predictive value. The validation results are consistent with the cross-validation findings. On both splits, all models that use bank transaction data outperform those based solely on application information. The combined features lead to the best result for all models. In general, these findings strongly support the hypothesis that bank statement transaction data offers significant predictive power in credit risk assessment for MSMEs in Malaysia. While application form information provides some discriminatory power, the inclusion of bank transaction-based features substantially enhances model performance, reinforcing the effectiveness of our proposed workflow.
7 Deployment and Operational Framework
The sixth phase of CRISP–DM focuses on the deployment of the proposed cash flow underwriting workflow in a production environment. This workflow is designed to leverage verifiable transaction data to enable transparent and evidence-based credit decisions while establishing a continuous feedback cycle for data collection, model retraining, and performance monitoring within existing core banking systems. To support this, a robust machine learning operations framework is implemented to structure the end-to-end pipeline into five key stages. Lastly, an integrated credit scoring framework is introduced to provide comprehensive risk evaluation for both established and new-to-credit MSMEs.
Data Ingestion
The pipeline begins when bank statements are received. Extraction modules parse and structure the raw transaction data using OCR and layout analysis techniques. This process transforms unstructured documents into a standardised tabular format suitable for downstream processing. The ingestion stage also performs integrity verification and data validation checks to ensure document authenticity.
Feature Engineering and Feature Store
Following ingestion, structured data are fed into the feature engineering module, which computes transaction-derived and behavioural features. These features are time-stamped, versioned, and stored in a centralised repository. The feature store provides a single source of truth for both training and inference, ensuring the same logic used to generate historical features is consistently applied to new applicants in real time. This design guarantees reproducibility, prevents feature skew, and supports traceability across model iterations.
Continuous Integration (CI)
The CI pipeline automates model retraining and validation prior to integration into the production registry. It is triggered by two primary events: (1) a predefined schedule (e.g., annually) when new labelled data become available from the core banking system, and (2) alerts from the monitoring system indicating potential model or data drift. Upon activation, the pipeline retrieves the latest versioned feature set and ground truth labels, retrains the model, and evaluates its performance against the current production model using predefined and consistent evaluation metrics.
Model Registry and Continuous Deployment (CD)
If the retrained model demonstrates superior performance, it is automatically versioned and stored in the model registry. The CD process then packages the validated model into a containerised microservice and deploys it as a secure REST API endpoint. To ensure reliability and minimise production risk, a canary deployment strategy is applied—initially routing a small fraction of live traffic to the new model for performance validation before full-scale rollout.
Continuous Monitoring and Feedback Loop
Once the bank statement-based credit scoring model is deployed, it operates within a Champion–Challenger framework (kim2019champion) to enable continuous performance optimisation. The Champion model serves as the current production baseline, while one or more Challenger models are periodically retrained on the latest data and evaluated in parallel. Model performance is continuously monitored and compared across predefined metrics. When a Challenger consistently outperforms the Champion, the system automatically notifies designated bank officers to review the results and promote the new model to production. This approach ensures sustained model improvement, robustness, and operational stability to maintain optimal performance.
Integrated Credit Scoring Framework
In production, the cash flow underwriting system functions as an integrated decision engine embedded within the lending process. Upon submission of bank statements, AI modules automatically extract and structure the raw data using OCR and natural language processing models. The processed data then undergoes fraud detection, cash flow analysis, and network analytics to generate transaction-derived features and credit scores for MSME applicants. The existing bureau-based scorecard, designed for borrowers with established credit histories, operates in parallel with the newly developed cash flow–based model. Each model independently produces a risk rating, and a risk override mechanism ensures conservative decisioning; if either model indicates higher risk, the final classification adopts that rating. This framework provides a transparent, evidence-based, and data-driven approach to MSME credit assessment using adaptive, explainable AI scoring. It also extends financial access to new-to-credit businesses through verifiable transaction data.
8 Conclusion
This study is the first to introduce bank statement transactions as alternative data for MSME loan underwriting in Malaysia. Our study demonstrates that transaction-derived features capture dynamic aspects of MSMEs’ financial behaviour that are often overlooked in traditional credit models. These features also exhibit strong predictive power with high IV scores. Empirical results show that models built solely on application information perform modestly, whereas those based on bank transaction data achieve substantially better results. The combined features further improve the model’s predictive power, confirming the complementary nature and practical value of alternative data.
9 Acknowledgments
We thank the Cradle Fund and the Microsoft for Startups program for their continuous support. This work was made possible through the collective contributions of the DocLab, DocSuite and MLSuite teams at AI Lens, whose efforts were instrumental in the development of this research.
10 Limitations and Future Work
This work has several limitations that should be addressed in future research. First, the dataset comprises 611 loan applications from a single Malaysian lending institution. While the dataset represents real-world production data collected over a two-year period, the sample size is constrained by the practical realities of MSME lending in emerging markets, where data availability remains limited due to stringent data privacy and regulatory requirements. Future work should validate these findings across multiple institutions and larger datasets as additional MSME lending data become available.
Second, the class imbalance in our dataset (518 non-default vs. 93 default cases) reflects the natural distribution observed in real-world lending portfolios, where the majority of borrowers successfully repay their loans. While this imbalance poses modelling challenges, it accurately reflects real-world credit risk settings. In this work, we addressed this limitation through appropriate model selection, WOE-based feature engineering, and careful evaluation using AUROC, which is robust to class imbalance. It is important to note that class imbalance is not a limitation that can be fully eliminated, but rather an inherent characteristic of credit scoring datasets. As credit assessment models improve and lending decisions become more accurate, successful repayment rates increase, thereby maintaining or even intensifying this natural imbalance. Nevertheless, future studies could explore advanced resampling techniques or cost-sensitive learning methods to further enhance model performance on minority class prediction.
Third, while we describe multiple AI modules for document processing, fraud detection, and transaction analysis within the proposed end-to-end cash flow underwriting workflow, the evaluation focuses on overall credit scoring performance rather than the effectiveness of individual modules. However, detailed module-level assessment is constrained by the use of proprietary methods and sensitive operational data that cannot be publicly disclosed. Future work utilising synthetic or anonymised data could enable more granular module-level analysis and benchmarking. Additionally, future research could explore formalising these modules as fully autonomous agents, enabling the development of a multi-agent credit scoring architecture with clearer inter-agent responsibilities and interactions.
Fourth, the proposed workflow has been deployed in a production environment with one Malaysian lending institution. While this demonstrates real-world applicability, broader validation across different lending institutions, regulatory environments, and MSME segments would strengthen the generalisability of our findings. Further research is needed to systematically integrate the proposed cash flow underwriting workflow with existing credit assessment systems that rely on behavioural and credit bureau data, ensuring coherent and robust risk evaluation. Additionally, longitudinal studies examining model performance stability across economic cycles would provide valuable insights into the temporal robustness and adaptability of bank statement-based features under varying macroeconomic conditions.